# Hilbert space fragmentation from lattice geometry

### v1

@@ Abstract
The eigenstate thermalization hypothesis describes how isolated many-body quantum systems reach thermal equilibrium. However, quantum many-body scars and Hilbert space fragmentation violate this hypothesis and cause nonthermal behavior. We demonstrate that Hilbert space fragmentation may arise from lattice geometry in a spin-$1/2$ model that conserves the number of domain walls. We generalize a known, one-dimensional, scarred model to larger dimensions and show that this model displays Hilbert space fragmentation on the Vicsek fractal lattice of arbitrary generation. Using Monte Carlo methods, the model is characterized as strongly fragmented on the Vicsek fractal lattice when the number of domain walls is either small or close to the maximal value. We show that the model displays signatures similar to Hilbert space fragmentation on the second-generation hexaflake fractal lattice, the two-dimensional lattice, and a modified two-dimensional lattice. We study the autocorrelation function of local observables and demonstrate that the model displays nonthermal dynamics.
@@ I Introduction
The nature of fragmented models depends on the system dimension. For instance, the $t$$J_{z}$ model displays HSF in one dimension [26], but it is not fragmented in two dimensions [27]. Likewise, while the pair-flip model displays HSF in one dimension, it is scarred in dimensions larger than one [49, 27]. Hence, the presence of HSF in one dimension does not guarantee its existence in higher dimensions. Several works have, however, constructed models that are HSF in dimensions larger than one [34, 50, 51]. These findings call for a better understanding of the connection between HSF and the number of spatial dimensions. Besides analyzing fragmentation in integer dimensions, a general analysis would naturally study HSF on finite generation fractal lattices. Fractal lattices have a more complicated structure, and one may wonder whether HSF can arise from the lattice geometry itself. If this is possible, then lattices may be constructed where HSF is present on certain sublattices but not on other sublattices.
@@
In this work, we generalize a known, domain-wall conserving, one-dimensional, scarred model [17] to lattices of dimension larger than one. While the one-dimensional model hosts a few nonthermal scar states, the Hilbert space shatters into numerous dynamically disconnected subspaces when the model is placed in higher dimensions. We study the model on the Vicsek fractal lattice, hexaflake fractal lattice, two-dimensional lattice, and a modified two-dimensional lattice. We show that the model displays HSF on the Vicsek fractal lattice of any generation and that the fragmentation is strong in symmetry sectors where the number of domain walls is either close to zero or close to the maximal number of domain walls. We observe features similar to HSF on all other considered lattices and we find that the level of fragmentation depends on the lattice geometry. This result is explained by lattice sites with more than two nearest neighbors confining the movement of domain walls. The combination of domain wall conservation and conservation of total magnetization was previously shown to cause HSF in a one-dimensional system [35], but we show that the domain wall conservation can be sufficient in higher dimensions. We demonstrate the nonthermal nature of the model by studying the time-averaged autocorrelation function of a local observable, and we compare the results with the Mazur bound.
@@
In Sec. II, we take a one-dimensional, domain-wall conserving, scarred model as our starting point and generalize the model to larger dimensional lattices. We present the Vicsek fractal lattice, hexaflake fractal lattice, two-dimensional lattice, and a modified two-dimensional lattice. In Sec. III, we demonstrate that the model displays HSF on the Vicsek fractal lattice of arbitrary generation by explicitly constructing an exponential number of Krylov subspaces. Furthermore, we observe that the model displays characteristics similar to HSF on a section of the second-generation hexaflake fractal lattice, the two-dimensional lattice, and a modified two-dimensional lattice. We describe the mechanisms that restrict the movement of domain walls and cause fragmentation. For the Vicsek fractal lattice, we estimate the dimensions of symmetry sectors with a small number of domain walls using Monte Carlo importance sampling. We also estimate the dimension of the largest Krylov subspace in each symmetry sector. These results show that the largest Krylov subspace represents a vanishingly small part of the full symmetry sector. Hence, the model is strongly fragmented on the Vicsek fractal lattice for a small number of domain walls. We extend this result to symmetry sectors where the number of domain walls is close to the maximal value. In Sec. IV, we study the long-time average of the autocorrelation function of a local observable and compare the results with the Mazur bound. While the time-averaged autocorrelation function does not converge to the Mazur bound for the considered lattices, the Mazur bound becomes tight when the Hamiltonian is perturbed by a block diagonal random matrix. In Sec. V, we summarize the results.
@@ II Model
Figure 1: Illustration of the considered lattices. Black dots show the lattice sites and red lines display nearest neighbor edges. (a) The first-generation Vicsek fractal lattice. (b) The second-generation Vicsek fractal lattice is obtained from generation one by substituting all lattice sites with five new sites. (c) Similarly, the third-generation Vicsek fractal lattice is obtained from generation two by substituting all sites with generation one Vicsek fractal lattices. (d) Second-generation hexaflake fractal lattice. We consider the lattice consisting of three connected first-generation hexaflake fractal lattices (above the dashed line). (e) Two-dimensional lattice of size $5\times 4$. (f) Modified two-dimensional lattice constructed from four connected first-generation Vicsek fractal lattices.
@@ III Hilbert space fragmentation
where $\mathcal{H}$ is the full Hilbert space, $\mathcal{K}_{i}$ denotes a Krylov subspace of dimension $d_{i}$, and $\lvert\psi_{i}\rangle$ is a product state that generates $\mathcal{K}_{i}$.
@@
We study the fragmentation of the model in Eq. (2) on different lattices by computing the size and number of Krylov subspaces. Let $|s_{1}\ldots s_{N}\rangle$ be a simultaneous eigenket of the Pauli matrices in the $z$-direction $\{\sigma_{\bm{r}}^{z}\}$. We compute the Krylov subspace generated by a state $\lvert s_{1}\ldots s_{N}\rangle$ by iterative look-up in the matrix representation of the kinetic part of the Hamiltonian $H_{\lambda}$. In essence, if $\langle s_{1}^{\prime}\ldots s_{N}^{\prime}|H_{\lambda}|s_{1}\ldots s_{N}\rangle\neq 0$ then $|s_{1}^{\prime}\ldots s_{N}^{\prime}\rangle$ resides to the same Krylov subspace as $|s_{1}\ldots s_{N}\rangle$. We utilize the conservation of the number of domain walls by considering each symmetry sector separately. For all considered lattices, we generally observe the Hilbert space separating into Krylov subspaces. However, the degree of fragmentation depends on the number of domain walls and the geometry of the lattice. Figure 3 displays the matrix representation of the Hamiltonian operator in the basis $\{|s_{1}\ldots s_{N}\rangle\}$ for each lattice. We show each symmetry sector separately, and the basis states are arranged to allow the block diagonal structure from Eq. (III). The gray pixels represent nonzero matrix elements while white pixels represent vanishing matrix elements.
@@
Figure 3: Depictions of the Hamiltonian $H$ for each symmetry sector. Gray pixels represent nonzero matrix elements and white pixels correspond to vanishing matrix elements. The figure illustrates the block diagonal structure of the Hamiltonian operator due to Hilbert space fragmentation. $n_{\text{dw}}$ denotes the number of domain walls characterizing the symmetry sector, and $D\equiv D_{n_{\text{dw}}}$ denotes the dimension of the symmetry sector. The red asterisks mark symmetry sectors containing just a single Krylov subspace that spans the full sector. For the Vicsek fractal lattice, the Hamiltonian is only depicted up to $n_{\text{dw}}=18$ domain walls since the Krylov subspaces in the symmetry sector with $n_{\text{dw}}$ domain walls have the same sizes as the Krylov subspaces in the symmetry sector with $n_{\text{dw}}^{\text{max}}-n_{\text{dw}}$ domain walls as discussed in Appendix A.
III.1 Vicsek fractal lattice
@@
The kinetic term in the Hamiltonian operator $H_{\lambda}$ on the Vicsek fractal lattice is invariant under the transformation that inserts domain walls on edges where there is no domain wall and removes domain walls from edges where there is a domain wall. Therefore, the size and number of Krylov subspaces in the sector with $n_{\text{dw}}$ domain walls are identical to size and number of the Krylov subspaces in the sector with $n_{\text{dw}}^{\text{max}}-n_{\text{dw}}$ domain walls, where $n_{\text{dw}}^{\text{max}}$ is the maximal number of domain walls on the Vicsek fractal lattice of generation $g$. Without loss of generality, we study sectors with less than or equal to half filling of domain walls, i.e., $n_{\text{dw}}\leq n_{\text{dw}}^{\text{max}}/2$. We further discuss this point in Appendix A.
@@
Figure 3(a) displays the Hamiltonian operator for the second-generation Vicsek fractal lattice. We observe fragmentation within all symmetry sectors, with the amount of fragmentation varying between sectors. The fragmentation is caused by lattice sites with more than two nearest neighbors which may restrict the movement of domain walls. For instance, if all domain walls are located within one “arm” of the Vicsek fractal lattice, then the central spin can not be flipped, and it will act as a blockade for the movement of the domain walls as illustrated in Fig. 2. The dynamics of each Krylov subspace is, therefore, restricted to isolated regions on the lattice, with the remaining sites being frozen. In Appendix B, we formally demonstrate this point. We prove that the model displays HSF on the Vicsek fractal lattice of any generation by explicitly constructing an exponential number of one-dimensional Krylov subspaces.
@@
III.2 The hexaflake fractal lattice and two-dimensional lattice
Recall that sites with four or more nearest neighbors can restrict the movement of domain walls by acting as blockades. The second-generation hexaflake fractal lattice and the two-dimensional lattice are characterized by all sites having four or six nearest neighbors. Hence, we expect the model to display a large degree of fragmentation on these lattices since every site may act as a blockade. Figures 3(b) and 3(c) display the symmetry sectors of the Hamiltonian operator. As expected, the Hamiltonian operators on both lattices are split into numerous dynamically disconnected subspaces for a small number of domain walls similar to the result for the Vicsek fractal. However, for a large number of domain walls, the symmetry sectors display almost no fragmentation. We interpret this result as sites failing to act as blockades when the number of domain walls is large enough.
III.3 The modified two-dimensional lattice

### v2

@@ Abstract
The eigenstate thermalization hypothesis describes how isolated many-body quantum systems reach thermal equilibrium. However, quantum many-body scars and Hilbert space fragmentation violate this hypothesis and cause nonthermal behavior. We demonstrate that Hilbert space fragmentation may arise from lattice geometry in a spin-$1/2$ model that conserves the number of domain walls. We generalize a known, one-dimensional, scarred model to larger dimensions and show that this model displays Hilbert space fragmentation on the Vicsek fractal lattice and the two-dimensional lattice. Using Monte Carlo methods, the model is characterized as strongly fragmented on the Vicsek fractal lattice when the number of domain walls is either small or close to the maximal value. On the two-dimensional lattice, the model is strongly fragmented when the density of domain walls is low and weakly fragmented when the density of domain walls is high. Furthermore, we show that the fragmentation persists at a finite density of domain walls in the thermodynamic limit for the Vicsek fractal lattice and the two-dimensional lattice. We also show that the model displays signatures similar to Hilbert space fragmentation on a section of the second-generation hexaflake fractal lattice and a modified two-dimensional lattice. We study the autocorrelation function of local observables and demonstrate that the model displays nonthermal dynamics.
@@ I Introduction
The nature of fragmented models depends on the system dimension. For instance, the $t$$J_{z}$ model displays HSF in one dimension [26], but it is not fragmented in two dimensions [27]. Likewise, while the pair-flip model displays HSF in one dimension, it is scarred in dimensions larger than one [49, 27]. Hence, the presence of HSF in one dimension does not guarantee its existence in higher dimensions. Several works have, however, constructed models that are HSF in dimensions larger than one [34, 50, 51, 52, 53]. These findings call for a better understanding of the connection between HSF and the number of spatial dimensions. Besides analyzing fragmentation in integer dimensions, a general analysis would naturally study HSF on fractal lattices. Fractal lattices have a more complicated structure, and one may wonder whether HSF can arise from the lattice geometry itself. If this is possible, then lattices may be constructed where HSF is present on certain sublattices but not on other sublattices.
@@
In this work, we generalize a known, domain-wall conserving, one-dimensional, scarred model [17] to lattices of dimensions larger than one. While the one-dimensional model hosts a few nonthermal scar states, the Hilbert space shatters into numerous dynamically disconnected subspaces when the model is placed in higher dimensions. We study the model on the Vicsek fractal lattice, the two-dimensional lattice, a section of the second-generation hexaflake fractal lattice, and a modified two-dimensional lattice. We show that the model displays HSF on the Vicsek fractal lattice and the two-dimensional lattice. For the Vicsek fractal lattice, the fragmentation is strong in symmetry sectors where the number of domain walls is either close to zero or close to the maximal number of domain walls. For the two-dimensional lattice, the model is strongly fragmented when the density of domain walls is low and weakly fragmented when the density of domain walls is high. We also show that the fragmentation persists at a finite domain wall density in the thermodynamic limit on the Vicsek fractal lattice and the two-dimensional lattice. We observe features similar to HSF on a section of the second-generation hexaflake fractal lattice and the modified two-dimensional lattice. Furthermore, we find that the level of fragmentation depends on the lattice geometry. This result is explained by lattice sites with more than two nearest neighbors confining the movement of domain walls. The combination of domain wall conservation and conservation of total magnetization was previously shown to cause HSF in a one-dimensional system [35]. Furthermore, domain wall conservation has been shown to cause HSF on the two-dimensional lattice with periodic boundary conditions [50, 51]. We show that domain wall conservation also causes HSF on other lattices with dimensions larger than one. We demonstrate the nonthermal nature of the model by studying the time-averaged autocorrelation function of a local observable, and we compare the results with the Mazur bound.
@@
In Sec. II, we take a one-dimensional, domain-wall conserving, scarred model as our starting point and generalize the model to larger dimensional lattices. We present the Vicsek fractal lattice, the hexaflake fractal lattice, the two-dimensional lattice, and a modified two-dimensional lattice. In Sec. III, we demonstrate that the model displays HSF on the Vicsek fractal lattice and the two-dimensional lattice by explicitly constructing an exponential number of Krylov subspaces for both lattices. Furthermore, we observe that the model displays characteristics similar to HSF on a section of the second-generation hexaflake fractal lattice and a modified two-dimensional lattice. We describe the mechanisms that restrict the movement of domain walls and cause fragmentation. For the Vicsek fractal lattice, we estimate the dimensions of the symmetry sectors with a small number of domain walls using Monte Carlo importance sampling. We also estimate the dimension of the largest Krylov subspace in each symmetry sector. These results show that the largest Krylov subspace represents a vanishingly small part of the full symmetry sector. Hence, the model is strongly fragmented on the Vicsek fractal lattice for a small number of domain walls. We extend this result to symmetry sectors where the number of domain walls is close to the maximal value. For the two-dimensional lattice, we compute the dimensions of the Krylov subspaces and symmetry sectors exactly for various system sizes. Based on this data, we characterize the model as strongly fragmented when the density of domain walls is low and weakly fragmented when the density of domain walls is high. In Sec. IV, we study the long-time average of the autocorrelation function of a local observable and compare the results with the Mazur bound. While the time-averaged autocorrelation function does not converge to the Mazur bound for the considered lattices, the Mazur bound becomes tight when the Hamiltonian is perturbed by a block diagonal random matrix. In Sec. V, we summarize the results.
@@ II Model
Figure 1: Illustration of the considered lattices. Black dots show the lattice sites and red lines display nearest neighbor edges. (a) The first-generation Vicsek fractal lattice. (b) The second-generation Vicsek fractal lattice is obtained from generation one by substituting all lattice sites with five new sites. (c) Similarly, the third-generation Vicsek fractal lattice is obtained from generation two by substituting all sites with generation one Vicsek fractal lattices. (d) The second-generation hexaflake fractal lattice. We consider the lattice consisting of three connected first-generation hexaflake fractal lattices (above the dashed line). (e) The two-dimensional lattice of size $L_{x}\times L_{y}=5\times 4$. (f) The modified two-dimensional lattice constructed from four connected first-generation Vicsek fractal lattices.
@@ III Hilbert space fragmentation
where $\mathcal{H}$ is the full Hilbert space, $\mathcal{K}_{i}$ denotes a Krylov subspace of dimension $d_{i}$, and $\lvert\psi_{i}\rangle$ is a product state that generates $\mathcal{K}_{i}$. Let $|s_{1}\ldots s_{N}\rangle$ be a simultaneous eigenket of the spin operators in the $z$-direction, i.e., $\{\sigma_{\bm{r}}^{z}/2\}$ where $\sigma_{\bm{r}}^{z}$ is the Pauli $z$-operator on site $\bm{r}$. The Krylov subspace generated by a state $\lvert s_{1}\ldots s_{N}\rangle$ may be computed by iterative look-up in the matrix representation of the kinetic part of the Hamiltonian $H_{\lambda}$. In essence, if $\langle s_{1}^{\prime}\ldots s_{N}^{\prime}|H_{\lambda}|s_{1}\ldots s_{N}\rangle\neq 0$ then $|s_{1}^{\prime}\ldots s_{N}^{\prime}\rangle$ belong to the same Krylov subspace as $|s_{1}\ldots s_{N}\rangle$. Since the Hamiltonian operator conserves the number of domain walls, the Krylov subspaces may be constructed by considering each symmetry sector separately.
@@
In this section, we demonstrate that the model from Eq. (2) displays HSF on certain lattices and we characterize the fragmentation in various symmetry sectors. However, before delving into this analysis, it is instructive to inspect the fragmentation of the model for a single system size of the considered lattices. We consider the second-generation Vicsek fractal lattice with $25$ dynamically active sites from Fig. 1(b), the section of the second-generation hexaflake fractal lattice with $21$ dynamically active sites from Fig. 1(d), the two-dimensional lattice of size $L_{x}\times L_{y}=5\times 4$ from Fig. 1(e), and the modified two-dimensional lattice with $20$ dynamically active sites from Fig. 1(f). The system sizes are chosen such that the considered lattices consist of approximately the same number of sites. Figure 3 displays the matrix representation of the Hamiltonian operator in the basis $\{|s_{1}\ldots s_{N}\rangle\}$ for the single system size of each lattice. We show each symmetry sector separately, and the basis states are arranged to allow the block diagonal structure from Eq. (4). The gray pixels represent nonzero matrix elements while white pixels represent vanishing matrix elements. For all considered lattices, we generally observe that the Hilbert space shatters into numerous Krylov subspaces. However, the degree of fragmentation depends on the number of domain walls and the geometry of the lattice.
@@
Figure 3: The Hamiltonian in each symmetry sector for (a) the second-generation Vicsek fractal lattice illustrated in Fig. 1(b) with $25$ dynamically active sites, (b) the section of the second-generation hexaflake fractal lattice illustrated in Fig. 1(d) with $21$ dynamically active sites, (c) the two-dimensional lattice illustrated in Fig. 1(e) of size $L_{x}\times L_{y}=5\times 4$, and (d) the modified two-dimensional lattice illustrated in Fig. 1(f) with $20$ dynamically active sites. Gray pixels represent nonzero matrix elements and white pixels correspond to vanishing matrix elements. The figure illustrates the block diagonal structure of the Hamiltonian operator due to Hilbert space fragmentation. $n_{\text{dw}}$ denotes the number of domain walls characterizing the symmetry sector, and $D\equiv D_{n_{\text{dw}}}$ denotes the dimension of the symmetry sector. For the Vicsek fractal lattice, the Hamiltonian is only depicted up to $n_{\text{dw}}=18$ domain walls since the Krylov subspaces in the symmetry sector with $n_{\text{dw}}$ domain walls have the same sizes as the Krylov subspaces in the symmetry sector with $n_{\text{dw}}^{\text{max}}-n_{\text{dw}}$ domain walls as discussed in Appendix A. The red asterisks mark symmetry sectors containing just a single Krylov subspace that spans the full sector.
III.1 The Vicsek fractal lattice
@@
The kinetic term in the Hamiltonian operator $H_{\lambda}$ on the Vicsek fractal lattice is invariant under the transformation that inserts domain walls on edges where there is no domain wall and removes domain walls from edges where there is a domain wall. Therefore, the size and number of Krylov subspaces in the sector with $n_{\text{dw}}$ domain walls are identical to the size and number of the Krylov subspaces in the sector with $n_{\text{dw}}^{\text{max}}-n_{\text{dw}}$ domain walls, where $n_{\text{dw}}^{\text{max}}$ is the maximal number of domain walls on the Vicsek fractal lattice of generation $g$. Without loss of generality, we study sectors with less than or equal to half filling of domain walls, i.e., $n_{\text{dw}}\leq n_{\text{dw}}^{\text{max}}/2$. We further discuss this point in Appendix A.
@@
Figure 3(a) displays the Hamiltonian operator for the second-generation Vicsek fractal lattice consisting of $25$ dynamically active sites. We observe fragmentation within all symmetry sectors, with the amount of fragmentation varying between sectors. The fragmentation is caused by lattice sites with more than two nearest neighbors which may restrict the movement of domain walls. For instance, if all domain walls are located within one “arm” of the Vicsek fractal lattice, then the central spin can not be flipped, and it will act as a blockade for the movement of the domain walls as illustrated in Fig. 2. The dynamics of each Krylov subspace is, therefore, restricted to isolated regions on the lattice, with the remaining sites being frozen. Following this line of thought, we prove in Appendix B that the model displays HSF on the Vicsek fractal lattice. In the proof, we construct an exponential number of one-dimensional Krylov subspaces by trapping the domain walls in certain regions of the lattice. We also show that the fragmentation persists at finite domain wall densities in the thermodynamic limit.
@@
III.2 The two-dimensional lattice
Figure 5: (a) The number of Krylov subspaces as a function of system size $L_{x}\times L_{y}$. We vary the length of one side of the lattice $L_{x}$ while keeping the other side fixed at $L_{y}=2$ (dark blue dots), $L_{y}=3$ (light blue triangles), and $L_{y}=4$ (light red diamonds). We also increase the length $L$ of both sides simultaneously $L_{x}\times L_{y}=L\times L$ (dark red crosses). Notice that some data points are included in more than one of these groups and have multiple markers, e.g., $L_{x}\times L_{y}=2\times 2$. The number of Krylov subspaces scales exponentially with system size and is larger than the lower bound (dashed line) obtained in Appendix D. (b)-(d) the ratio between the dimension of the largest Krylov subspace and the dimension of the corresponding symmetry sector as a function of system size for various densities of domain walls. We consider (b) one-quarter filling $n_{\text{dw}}=n_{\text{dw}}^{\text{max}}/4$, (c) half-filling $n_{\text{dw}}=n_{\text{dw}}^{\text{max}}/2$, and (d) three-quarter filling $n_{\text{dw}}=3n_{\text{dw}}^{\text{max}}/4$. The results indicate that the model is strongly fragmented for a low density of domain walls and weakly fragmented for a high density of domain walls.
Next, we study the model from Eq. (2) on the two-dimensional lattice of size $L_{x}\times L_{y}=5\times 4$. Figure 3(c) displays the Hamiltonian operator in each symmetry sector as a block diagonal matrix. The figure shows that symmetry sectors with a small number of domain walls shatter into numerous Krylov subspaces and the figure hints that the model is fragmented on the two-dimensional lattice. We formally prove this statement in Appendix D by computing a lower bound on the number of Krylov subspaces. We show that this lower bound scales exponentially with system size, and hence, that the model displays HSF on the two-dimensional lattice. We also demonstrate that the fragmentation persists at a finite density of domain walls in the thermodynamic limit. Figure 5(a) shows the number of Krylov subspaces for various system sizes $L_{x}\times L_{y}$ and the lower bound from Appendix D. The figure illustrates that the number of Krylov subspaces scales exponentially with system size and that the number of Krylov subspaces grows faster with system size than the lower bound.
We characterize the fragmentation of the model on the two-dimensional lattice by computing the Krylov subspaces in certain symmetry sectors for different system sizes. Similarly to the analysis in Sec. III.1, we study the ratio between the dimension $d_{n_{\text{dw}}}^{\text{max}}$ of the largest Krylov subspace in the symmetry sector with $n_{\text{dw}}$ domain walls and the dimension $D_{n_{\text{dw}}}$ of the corresponding symmetry sector. The system size is varied by extending one side of the lattice while keeping the length of the other side fixed, e.g., the system sizes $L_{x}\times 4$ for different values of $L_{x}$. We also vary the system size by increasing the length of both sides simultaneously, i.e., the square lattices $L_{x}\times L_{y}=L\times L$ for different values of $L$. Figures 5(b)-(d) display the ratio $d_{n_{\text{dw}}}^{\text{max}}/D_{n_{\text{dw}}}$ as a function of system size. We consider symmetry sectors with a specific filling of domain walls, e.g., half-filling $n_{\text{dw}}=n_{\text{dw}}^{\text{max}}/2$, and we round to the nearest valid number of domain walls when necessary. For one-quarter filling $n_{\text{dw}}=n_{\text{dw}}^{\text{max}}/4$, we find that the ratio $d_{n_{\text{dw}}}^{\text{max}}/D_{n_{\text{dw}}}$ decreases with increasing system size both when increasing the length of one side and when increasing both sides simultaneously. This result indicates that the model displays strong fragmentation on the two-dimensional lattice for a low density of domain walls. At half-filling $n_{\text{dw}}=n_{\text{dw}}^{\text{max}}/2$, the ratio decreases for system sizes $L_{x}\times 2$ when increasing $L_{x}$. However, the ratio increases for the lattices $L_{x}\times 3$ and $L_{x}\times 4$ when increasing $L_{x}$ and for the square lattices $L\times L$ when increasing $L$. At three-quarter filling $n_{\text{dw}}=3n_{\text{dw}}^{\text{max}}/4$, the ratio increases and approaches unity when increasing the length of one side or both sides. This result indicates that the model is weakly fragmented at a high density of domain walls. Recall that sites with four or more nearest neighbors can restrict the movement of domain walls by acting as blockades. On the two-dimensional lattice, all sites have four nearest neighbors and every site may, therefore, limit the movement of the domain walls. Indeed, we find that the model is strongly fragmented in symmetry sectors with a low density of domain walls. However, we observe that the model is weakly fragmented in symmetry sectors with a high density of domain walls. We interpret this result as sites failing to act as blockades when the density of domain walls is sufficiently large.
We remark that a model similar to Eq. (2) was previously studied on the two-dimensional lattice with periodic boundary conditions [50, 51]. This model was proven to display HSF when both side lengths of the lattice are a multiple of an odd integer larger than one [50]. This argument also applies to our model, but for completeness, we provide the proof in Appendix D which is valid for all system sizes. The model with periodic boundary conditions was also shown to display strong fragmentation for a low density of domain walls and weak fragmentation for a high density of domain walls [51]. However, the presence of respectively weak and strong fragmentation for periodic boundary conditions does not guarantee the same fragmentation strength for open boundary conditions. Indeed, the domain walls are more confined for open boundary conditions and one might naively expect that the model will display stronger fragmentation in this case. Interestingly, we observe that the fragmentation displays the same behavior as a function of domain wall density for open boundary conditions as for periodic boundary conditions.
III.3 The hexaflake fractal lattice and the modified two-dimensional lattice
Finally, we consider the model on the section of the second-generation hexaflake fractal lattice with $21$ dynamically active sites from Fig. 1(d) and the modified two-dimensional lattice with $20$ dynamically active sites from Fig. 1(f). Figure 3(b) displays the Krylov subspaces in various symmetry sectors for the section of the second-generation hexaflake fractal lattice. All sites in this lattice have four or six nearest neighbors and, similarly to the two-dimensional lattice, every site may act as a blockade to the domain walls. We observe somewhat similar fragmentation on the section of the second-generation hexaflake fractal lattice as on the two-dimensional lattice of size $L_{x}\times L_{y}=5\times 4$. In particular, we observe that the model on the hexaflake fractal lattice is more fragmented in symmetry sectors with fewer domain walls as compared to symmetry sectors with many domain walls.